OpenStudy (anonymous):

FREE MEDAL!! Let f(x) = x2 - 16. Find f-1(x) would it be the sqrt of x+16????

OpenStudy (luigi0210):

This is where you would switch y and x and solve for y

OpenStudy (jdoe0001):

I take the medal, I mean, is free as in free drinks, so hey why not

OpenStudy (anonymous):

is it the sqrt of x+16?

OpenStudy (jdoe0001):

\(\bf f(x) \implies \color{red}{y} = \color{blue}{x}^2-16\\ f^{-1}(x) \implies \color{blue}{x} = \color{red}{y}^2-16\)

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

well \(\bf f^{-1}(x) \implies x = y^2-16 \implies x+16 = y^2\)

OpenStudy (luigi0210):

....

OpenStudy (jdoe0001):

so, you still need to take square root on both sides

OpenStudy (anonymous):

ok so that gives me the sqrt of x+16??????

OpenStudy (jdoe0001):

ohh yeah,.... shoot I didn't read you well, yes, it does \(\bf f^{-1}(x) \implies x = y^2-16 \implies x+16 = y^2 \implies \sqrt{ x+16 } = y\)